OpenStudy (alexh107):

Verify the trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation: cot x sec^4 x = cot x + 2 tan x + tan^3 x Please walk me through this step by step. I don't think I've ever been more lost in math.

OpenStudy (mathmale):

Are you sure there are two equal signs in this expression?

OpenStudy (alexh107):

Oh sorry, this is the correct one: cot x sec^4 x = cot x + 2 tan x + tan^3 x

OpenStudy (mathmale):

that makes a huge difference.

OpenStudy (mathmale):

I see you have no fewer than 3 trig functions in this equation. Can you think of a way to cut that to just 2 trig functions?

OpenStudy (alexh107):

Not really. The only thing I know is that you're supposed to start with the right side since it's more complex but after that I don't know what to do.

OpenStudy (mathmale):

Need to dig up a table of trig identiies here. Do you have one in front of y ou?

OpenStudy (alexh107):

I have the common ones written in my notes.

OpenStudy (mathmale):

which one involves (csc x)^2?

OpenStudy (alexh107):

1 + cot^2 = csc^2

OpenStudy (mathmale):

good. that's correct, accurate.

OpenStudy (mathmale):

I see that all terms in your equation involve either tan or cot, except for one: (sec x)^4. Agree or disagree?

OpenStudy (alexh107):

Agree

OpenStudy (mathmale):

So we might be able to simplify the identity by getting rid of that (sec x)^4. Any ideas on how to do that? How is the sec x function defined?

OpenStudy (alexh107):

I'm not really sure. Would we need to do something so that it's like (sec^2)(sec^2) since most of the identities involve sec^2 not sec^4?

OpenStudy (mathmale):

That's true. The secant function is defined as 1/ cos x. We don't want to introduce yet another trig function into this inequality, so should not use this. You are correct in that you can re-write (sec x)^4 as [(sec x)^2]^2. Agreed or not?

OpenStudy (alexh107):

Agreed

OpenStudy (mathmale):

Try using the identity for (sec x)^2 that we discussed earlier.

OpenStudy (alexh107):

I got (1 + tan^2) (1 + tan^2) since 1 + tan^2 = sec^2 and we have sec^4 but I'm not sure how to use the one I wrote earlier for 1 + cot^2 = csc^2

OpenStudy (mathmale):

Remember, we don't want to introduce or re-introduce more trig functions here. so (csc x)^2 is out. What is the relationship between (tan x)^2 and (sec x)^2?

OpenStudy (alexh107):

1 + tan^2 = sec^2

OpenStudy (mathmale):

If you can u se Equation Editor, let's use it. But your latest input is clear. enuf (altho you've left out the argument x.). Note that we have (sec x)^4 in that equation. How could we use your latest result to eliminate (sec x)^4 from the equation?

OpenStudy (mathmale):

this is Equation Editor output:\[\tan ^{2}x+1=\sec^2x\]

OpenStudy (mathmale):

Starting with this latest equation, find an equation for \[\sec^4x\]

OpenStudy (alexh107):

\[(\tan^2 x + 1)(\tan^2 x + 1) = \sec^4 x\]

OpenStudy (alexh107):

I don't think that's right but I'm not sure what to do.

OpenStudy (mathmale):

What you have is fine. this identity allows us to eliminate (sec x)^4 from the equation. Why not multiply out\[(\tan^2x+1)^2\]

OpenStudy (mathmale):

and replace (sec x)^4 with the result?

OpenStudy (alexh107):

So that would make the first side of the equation: \[\cot x (\tan^2 x +1)^2\]

OpenStudy (mathmale):

cot x sec^4 x = cot x + 2 tan x + tan^3 x or\[\cot x \sec^4 x = \cot x + 2 \tan x + \tan^3 x\] Yes, you have the left side correct. Notice that you also have cot x on the right side.

OpenStudy (mathmale):

Would it be possible to get rid of the cot x factor?

OpenStudy (mathmale):

As you showed me, you now have\[\cot x (\tan^2x + 2 \tan x + 1) \] on the left side, and this is getting to look a lot like the right side. Note that the cot and tan are reciprocals of one another; can you put that fact to good use?

OpenStudy (mathmale):

Is this the right side you have also? cot x + 2 tan x + tan^3 x\[\cot x + 2 \tan x + \tan^3 x\]

OpenStudy (alexh107):

Yes that is the right side. I'm just confused on the left how we went from \[\cot x (\tan^2 x +1)^2 \to \cot x + 2\tan x +1\]

OpenStudy (mathmale):

We've gone thru so much detail already, there's bound to be an oversight somewhere. But note that you can and should expand the square of (tan x)^2 + 1).

OpenStudy (mathmale):

I think you've learned a lot here and made some progress towards proving the identity. At this point I ask you which would benefit y ou more, to complete this proof or to move on to another problem.

OpenStudy (alexh107):

I can try to move on to another problem I suppose and maybe that one will be easier and I can come back to this one.

OpenStudy (mathmale):

I've copied this one down and am gtoing to work on it a bit more. If you'll promise to come back and review what we've done here, then I'd suggest you move on to another proof. Each of us has only so much time, and so it's essential to put every min. to good use.

OpenStudy (alexh107):

Okay. I found another one on my paper that looks slightly easier. I can go try to work on that for now.

OpenStudy (mathmale):

Hey, I think I've proven the one we were working on. I'll share that with you later. Post the new problem now...Ask a (new) question, not adding on to this already long discussion.

OpenStudy (alexh107):

Okay, the new problem I have is:\[1 + \sec^2 x \sin^2 x = \sec^2 x\]

OpenStudy (mathmale):

Definitely have finished the proof.

OpenStudy (alexh107):

That's great. Thank you for taking the time to help me.

OpenStudy (mathmale):

And you're supposed to prove this new identity?

OpenStudy (alexh107):

Yes

OpenStudy (mathmale):

1 + \sec^2 x \sin^2 x = \sec^2 x\[1 + \sec^2 x \sin^2 x = \sec^2 x\]

OpenStudy (mathmale):

the tan and sec functions pair up in an identity, as do the cot and csc.

OpenStudy (mathmale):

Also, the reciprocal of the cos is the sec function. Could you use either or both of these facts to begin simplifying the given equation?

OpenStudy (mathmale):

Note: the sine and cosine are most familiar to most of us, so y ou might benefit from eliminating the sec function in favor of its reciprocal, the cos.

OpenStudy (mathmale):

Simply take the original equation and substitute \[\frac{ 1 }{ \cos^2 x }\]

OpenStudy (mathmale):

for (sec x)^2.

OpenStudy (mathmale):

just on the left side.

OpenStudy (alexh107):

\[1 + 1/\cos^2x + \sin^2x\]

OpenStudy (mathmale):

Al, don't we haver multiplication in the left-hand term? You've introduced a 2nd addition.

OpenStudy (alexh107):

Oh my bad \[1+1/\cos^2x \sin^2x\]

OpenStudy (mathmale):

Basically right; need parentheses around the reciprocal of (cos x)^2. But anyway. What you have now is [1 / (cos x)^2\*[sin x]^2. OK? Can you simplify that?

OpenStudy (mathmale):

\[1+\frac{ \sin^2x }{ \cos^2x }=?\]

OpenStudy (mathmale):

Leave the 1 alone. The 2nd term is easily simplified.

OpenStudy (alexh107):

I'm struggling to find the identity to use. Would it be tan x = sin x/ cos x

OpenStudy (mathmale):

yes, preciselyl.

OpenStudy (alexh107):

So that means 1 + tan^2 = sec^2?

OpenStudy (mathmale):

Yes. is that true or not?

OpenStudy (alexh107):

True

OpenStudy (alexh107):

So that's the end of that proof?

OpenStudy (mathmale):

Then you've proven the identity.

OpenStudy (mathmale):

yes.

OpenStudy (mathmale):

Now, regarding the previous problem: We have \[\cot x(\tan^4x+2\tan^2x + 1) = \cot x + 2\tan x + \tan^3x.\]

OpenStudy (mathmale):

convert that cot x to 1 / tan x and then multiply everything within parentheses by 1 / tan x.

OpenStudy (alexh107):

\[2\tan x/\tan x + \tan^3 x /\tan x = \]

OpenStudy (mathmale):

Compare that to the right side. See how close you're getting?

OpenStudy (mathmale):

Where you should have 2 (tan x)^2, you have written 2 tan x. fix that please.

OpenStudy (alexh107):

\[2 (\tan x)^2/\tan x + \tan^3/ \tan x\]

OpenStudy (mathmale):

Simplify the left side, then again compare the left and right sides of y our equation.

OpenStudy (alexh107):

\[2 \tan x + \tan^2 x\]

OpenStudy (mathmale):

Aside from some very minor mistakes you've succeeded in proving this identiy. I encourage you to review our discussion back to what we typed in 8 minutes ago, and make the changes necessary to validate your proof. typing 2 tan x instead of 2 (tan x)^2 was one example. I so much appreciate your involving y ourself in this work as much as you have.

OpenStudy (mathmale):

As much as I'd like to continue, I need to shift my focus elsewhere. Hope to have the privilege of working wi th you again soon. :)

OpenStudy (alexh107):

I understand. Thank you for your help and spending so much time with me on this. I understood the other proof we completed and hopefully I will be able to figure this one out eventually.

OpenStudy (mathmale):

Cool. It's worth the effort! See you, Alex! Bye.